OpenStudy (anonymous):

If you have the following inequality: $\sup_{[x_{k-1},x_k]}f(x'-x_{k-1})+\sup_{[x',x_{k}]}f(x_k-x')\geq \sup_{[x_{k-1},x_k]}f(x_k-x_{k-1})$ Let A B C denote the respective supremums in order. What does my professor mean by saying A≤C & B≤C implies $A(x'-x_{k-1})+B(x_k-x')\leq C(x'-x_{k-1})+C(x_k-x')=C(x_k-x_{k-1})$

OpenStudy (swolesammy):

.

OpenStudy (anonymous):

God damn texing is a nightmare on this site.

OpenStudy (anonymous):

http://mathb.in/44622

OpenStudy (anonymous):

If you have the following inequality: $\sup_{[x_{k-1},x_k]}f(x'-x_{k-1})+\sup_{[x',x_{k}]}f(x_k-x')\geq \sup_{[x_{k-1},x_k]}f(x_k-x_{k-1})$ Let A B C denote the respective supremums in order. What does my professor mean by saying A≤C & B≤C implies $A(x'-x_{k-1})+B(x_k-x')\leq C(x'-x_{k-1})+C(x_k-x')=C(x_k-x_{k-1})$